Here's my question: Adult tickets for a play cost $11 and child tickets cost $10. If there were 23 people at a performance and the theater collected $246 from ticket sales, how many children attended the play? I got 16 children, but I have a feeling that's not right.
Answer by scott8148(6628)

16 is the number of adults (NOT children) __ was it a lucky guess?

A+C=23 __ A=23-C

11A+10C=246 __ substituting __ 11(23-C)+10C=246 __ 253-11C+10C=246 __ -C=-7